# A Note on a Recent Attempt to Prove the Irrationality of $\zeta(5)$

### v2

@@ 1. Introduction
In 1882, Lindemann proved that the number $\pi$ is transcendental (see, for example, [6, Theorem 1.2]). As a consequence, the even values $\zeta(2n)$ are irrational. In 1979, Apéry [1] famously demonstrated that the number $\zeta(3)$ is irrational. Subsequently, Beukers [3] provided an elementary proof of $\zeta(3)$ based on integral representations. In 2001, Keith Ball and Tanguy Rivoal [2, Théorème 1] applied Nesterenko’s criterion to prove that the $\mathbb{Q}$-vector space spanned by odd zeta values is infinite-dimensional:
@@
is irrational. Incidentally, the proof of the irrationality of $\zeta(3)$ can also be found in Zudilin’s recent book [8, Section 7].
@@
This note is organized as follows. In Section 2, we identify the logical flaws in the proof presented in [5]. In Section 3, we outline a standard approach to proving the irrationality of a number and highlight that the irrationality of $\zeta(2m+1)$ for general odd integers $12m+$ 1 remains an open and intriguing area of research.
@@ 2. Logical Fallacies in Suman’s Proof
In this section, we analyze the flaws in Shekhar Suman’s proof of [5, Theorem 1]. While the preliminary result, [5, Lemma 1], is correct and well-established, the proof of [5, Theorem 1] contains a critical logical error.
Suman claims that the linear Diophantine equation [5, Eq. (47)] has no integer solution, which he uses to derive the irrationality of $\zeta(5)$. However, we show that [5, Eq. (47)] indeed admits integer solutions, invalidating his conclusion.
@@
Here lies the critical error: the claim that equation [5, Eq. (48)] has no integer solutions is logically independent of whether $\zeta(5)$ is an integer or not. The well-established fact that $\zeta(5)$ is not an integer cannot be used to prove that [5, Eq. (48)] has no integer solutions. The existence of integer solutions for [5, Eq. (48)] at $n=1$ directly invalidates the induction base case and, consequently, the entire proof.
To summarize, the error in [5, Theorem 1] lies in conflating two unrelated facts: the solvability of the Diophantine equation [5, Eq. (48)] and the irrationality of $\zeta(5)$. The former is a purely algebraic property of the equation, while the latter is a number-theoretic property of the zeta function. By assuming $\zeta(5)$ is not an integer to argue against the solvability of [5, Eq. (48)], Suman undermines the logical foundation of his proof. As a result, Theorem 1 in [5] is incorrect.
In addition to the errors in the proof of [5, Theorem 1], similar logical fallacies can be identified in the proof of [5, Theorem 2], which aims to establish the irrationality of $\zeta(2m+1)$ for integers $m\geq 2$. While the specific details of the argument in [5, Theorem 2] differ from [5, Theorem 1], the underlying issue remains the same: the reliance on assumptions or conclusions that are independent of the solvability of the corresponding Diophantine equations.
@@ 3. A Standard Approach to Proving the Irrationality of a Number
Numerical verification shows that the $I_{n}$ in [5, Lemma 1] does not satisfy the condition (3) of Proposition 3.1. Therefore, $I_{n}$ cannot be used to prove the irrationality of $\zeta(5)$ via Proposition 3.1. In fact, mathematicians have been searching for a suitable integral representation of the form $a_{n}+b_{n}\zeta(5)$ that converges to 0, but such a representation has not yet been found. Therefore, the irrationality of $\zeta(2m+1)$ for $m\geq 2$ still remains an open problem worth exploring. Finally, finding a good integral representation for $\zeta(5)$ or even for $\zeta(2m+1)$ that can be used to approximate them continues to be a worthwhile pursuit.

### v3

@@ 1. Introduction
In 1882, Lindemann proved that the number $\pi$ is transcendental (see, for example, [7, Theorem 1.2]). As a consequence, the even values $\zeta(2n)$ are irrational. In 1979, Apéry [1] famously demonstrated that the number $\zeta(3)$ is irrational. Subsequently, Beukers [3] provided an elementary proof of $\zeta(3)$ based on integral representations. In 2001, Keith Ball and Tanguy Rivoal [2, Théorème 1] applied Nesterenko’s criterion to prove that the $\mathbb{Q}$-vector space spanned by odd zeta values is infinite-dimensional:
@@
is irrational. Incidentally, the proof of the irrationality of $\zeta(3)$ can also be found in Zudilin’s recent book [9, Section 7].
@@
This note is organized as follows. In Section 2, we identify the logical flaws in the proof presented in [6]. In Section 3, we outline a standard approach to proving the irrationality of a number and highlight that the irrationality of $\zeta(2m+1)$ for general odd integers $12m+$ 1 remains an open and intriguing area of research. In Section 4, we discuss recent advancements, including the groundbreaking work of Calegari, Dimitrov, and Tang [4], which introduces a powerful new method for proving irrationality and offers promising directions for future research.
@@ 2. Logical Fallacies in Suman’s Proof
In this section, we analyze the flaws in Shekhar Suman’s proof of [6, Theorem 1]. While the preliminary result, [6, Lemma 1], is correct and well-established, the proof of [6, Theorem 1] contains a critical logical error.
Suman claims that the linear Diophantine equation [6, Eq. (47)] has no integer solution, which he uses to derive the irrationality of $\zeta(5)$. However, we show that [6, Eq. (47)] indeed admits integer solutions, invalidating his conclusion.
@@
Here lies the critical error: the claim that equation [6, Eq. (48)] has no integer solutions is logically independent of whether $\zeta(5)$ is an integer or not. The well-established fact that $\zeta(5)$ is not an integer cannot be used to prove that [6, Eq. (48)] has no integer solutions. The existence of integer solutions for [6, Eq. (48)] at $n=1$ directly invalidates the induction base case and, consequently, the entire proof.
To summarize, the error in [6, Theorem 1] lies in conflating two unrelated facts: the solvability of the Diophantine equation [6, Eq. (48)] and the irrationality of $\zeta(5)$. The former is a purely algebraic property of the equation, while the latter is a number-theoretic property of the zeta function. By assuming $\zeta(5)$ is not an integer to argue against the solvability of [6, Eq. (48)], Suman undermines the logical foundation of his proof. As a result, Theorem 1 in [6] is incorrect.
In addition to the errors in the proof of [6, Theorem 1], similar logical fallacies can be identified in the proof of [6, Theorem 2], which aims to establish the irrationality of $\zeta(2m+1)$ for integers $m\geq 2$. While the specific details of the argument in [6, Theorem 2] differ from [6, Theorem 1], the underlying issue remains the same: the reliance on assumptions or conclusions that are independent of the solvability of the corresponding Diophantine equations.
@@ 3. A Standard Approach to Proving the Irrationality of a Number
Numerical verification shows that the $I_{n}$ in [6, Lemma 1] does not satisfy the condition (3) of Proposition 3.1. Therefore, $I_{n}$ cannot be used to prove the irrationality of $\zeta(5)$ via Proposition 3.1. In fact, mathematicians have been searching for a suitable integral representation of the form $a_{n}+b_{n}\zeta(5)$ that converges to 0, but such a representation has not yet been found. Therefore, the irrationality of $\zeta(2m+1)$ for $m\geq 2$ still remains an open problem worth exploring. Finally, finding a good integral representation for $\zeta(5)$ or even for $\zeta(2m+1)$ that can be used to approximate them continues to be a worthwhile pursuit.
4. Future Perspectives
Recently, Calegari, Dimitrov, and Tang [4] made significant progress in the study of irrationality and $\mathbb{Q}$-linear independence of special values in number theory. Their work established several groundbreaking results, including the irrationality of the classical Dirichlet $L$-value
$L(2,\chi_{-3})=\sum_{n=0}^{\infty}\left(\frac{1}{(3n+1)^{2}}-\frac{1}{(3n+2)^{2}}\right),$
and demonstrated its linear independence with $1$ and $\pi^{2}$ over $\mathbb{Q}$. This result, formalized as [4, Theorem A] in their paper, represents a significant advancement in our understanding of special values of $L$-functions. Furthermore, their techniques generalized Zagier’s constructions and provided a new arithmetic holonomy bound framework.
Their work also yielded [4, Corollary B], which proves the irrationality of numerous related series. For instance, they showed that certain linear combinations of $L(2,\chi_{-3})$, $\pi^{2}$ are irrational. These results expand the set of known irrational numbers derived from $L$-functions and highlight the potential of their methods for exploring similar problems.
Moreover, [4, Theorem C] extends their analysis to logarithmic values, establishing the irrationality of $\log\left(1+\frac{1}{m}\right)\log\left(1+\frac{1}{n}\right)$ under the condition $|\frac{m}{n}-1|<\frac{1}{10^{6}}$ and proving the $\mathbb{Q}$-linear independence of the following four expressions:
$1,\quad\log\left(1+\frac{1}{m}\right),\quad\log\left(1+\frac{1}{n}\right),\quad\log\left(1+\frac{1}{m}\right)\log\left(1+\frac{1}{n}\right).$
These findings represent a major breakthrough in the study of special values of zeta and $L$-functions, with implications for broader problems in number theory, transcendence theory, and the study of modular forms and related automorphic objects. The techniques introduced by Calegari, Dimitrov, and Tang provide a robust foundation for future exploration, including the study of other special values and related constants in mathematics.